OpenStudy (anonymous):

f(x) = x2 − 13 and g(x) = x + 5. What is the value of f[g(−4)]?

OpenStudy (anonymous):

@qam2k @whpalmer4 @wyattp17 @wtbryant @Whiplash1994 @wakoff @e.cociuba @razor99

OpenStudy (anonymous):

@xoya

OpenStudy (ivettef365):

first substitute the x for -4 on the g(x)

OpenStudy (ivettef365):

what do you get?

OpenStudy (anonymous):

f(-4)=(-4)2-13(-4)=(-4)+5

OpenStudy (ivettef365):

it was only on g(x)

OpenStudy (anonymous):

? \

OpenStudy (ivettef365):

ok, let start again, substitute on g(x) the x for -4, so you have g(x) = x + 5 g(x) = -4 + 5 so g(x) = 1

OpenStudy (ivettef365):

now that you have g(x) =1, then substitute the 1 for x on f(x)

OpenStudy (anonymous):

ok I did that

OpenStudy (ivettef365):

so what answer did you get

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

-21?

OpenStudy (anonymous):

nope

OpenStudy (ivettef365):

f(x) = x^2 - 13 f(x) = (1)^2 -13 = ???

OpenStudy (anonymous):

-12?

OpenStudy (ivettef365):

good job, so the value of f[g(-4)] = -12

OpenStudy (anonymous):

ty

OpenStudy (ivettef365):

yw :)